40m
QIL
Cryo_Lab
CTN
SUS_Lab
CAML
OMC_Lab
CRIME_Lab
FEA
ENG_Labs
OptContFac
Mariner
WBEEShop
|
OMC elog |
Not logged in |
 |
|
Mon Jul 18 14:20:09 2022, Koji, General, General, OMC #002 Plan Portal
|
Tue Jul 19 03:17:56 2022, Koji, General, General, OMC #002 Power Budget before mirror cleaning
|
Thu Jul 21 02:55:06 2022, Koji, General, General, OMC #002 Power Budget after mirror cleaning
|
Thu Jul 21 14:35:35 2022, Koji, General, General, Bond reinforcement blocks for the invar brackets 
|
Thu Jul 21 14:55:48 2022, Koji, General, General, Optics bonding for the backscatter measurement
|
Thu Jul 21 17:47:00 2022, Koji, General, General, The profile of the beam incident on the fiber input coupler
|
Fri Jul 22 00:31:17 2022, Koji, General, General, OMC #002 backscatter measurement
|
Fri Jul 22 17:47:38 2022, Koji, General, General, C&B request for the reinforcement blocks
|
Tue Jul 26 00:12:58 2022, Koji, General, General, High QE PD: QE measurements
|
Wed Jul 27 10:34:09 2022, Koji, General, General, High QE PD: QE measurements 2 
|
Wed Jul 27 10:09:51 2022, Koji, General, General, 4+4 wire clamp in hand
|
Wed Jul 27 10:34:30 2022, Koji, General, General, OMC #002 Protective FirstContact Paint
|
Wed Jul 27 23:52:18 2022, Koji, General, General, OMC #002 Cable bracket replacement (1)    
|
Thu Jul 28 00:28:15 2022, Koji, General, General, OMC #002 Cable bracket replacement (2)  
|
Thu Jul 28 00:46:47 2022, Koji, General, General, Subject: OMC #002 Cable bracket replacement (3)  
|
Wed Aug 10 18:42:27 2022, Koji, General, General, OMC #002 Cable bracket replacement (4)  
|
Thu Aug 11 15:24:57 2022, Koji, General, General, OMC #002 Cable bracket replacement (5)
|
Mon Aug 15 21:31:56 2022, Koji, General, General, OMC #002 Cable bracket replacement (6)  
|
Mon Aug 15 22:06:18 2022, Koji, General, General, OMC #002 new cable tie installed 
|
Mon Aug 15 22:43:35 2022, Koji, General, General, OMC #002 Delamination repair Part1 (1) 
|
Mon Aug 15 22:49:03 2022, Koji, General, General, OMC #002 Delamination repair Part1 (2) 7x
|
Wed Aug 24 02:51:23 2022, Koji, General, General, OMC #002 Delamination repair Part1 (3) 8x
|
Wed Aug 24 02:53:46 2022, Koji, General, General, OMC #002 Delamination repair Part2 (1)    
|
Wed Aug 24 03:20:59 2022, Koji, General, General, OMC #002 Delamination repair Part2 (2) 11x
|
Wed Aug 24 03:26:43 2022, Koji, General, General, OMC #002 Delamination repair Part2 (3) 8x
|
Wed Aug 24 02:57:43 2022, Koji, General, General, EP30-2 bonding setup  
|
Wed Aug 24 11:29:47 2022, Koji, General, General, OMC #002 ready for shipment    
|
Thu Aug 25 14:22:08 2022, Koji, General, General, LLO OMC #001 Ballast Mass investigation    
|
Thu Aug 25 20:05:00 2022, Koji, General, General, LLO OMC #001 Ballast Mass investigation
|
Fri Aug 26 22:29:02 2022, Koji, General, General, Tool Shipping Prep  
|
|
|
== Initial Preparation ==
- [Done] OMC #002 placement
- [Done] OMC #002 locking
- Details OMC ELOG 414
== Measurements ==
- [Done] Transmission / Power budget before FirstContact OMC ELOG 416
- [Done] Transmission / Power budget after FirstContact OMC ELOG 417
- [Done] Backscatter measurement with a new deflection optics
- [Done] Optics bonding done waiting for cure OMC ELOG 420 -> Returned the bond to Madeline OMC ELOG 424
- [Done] Backscatter measurement
- Measurement: 0.6 ppm OMC ELOG 422
- (Transmission is 10~60mW. If the backscatter is the order of 1ppm or less, we expect the light level is ~10nW. Can we really detect it? How? ... OK... last time the measurement has been done with the stick PD type powermeter with baffles and the room light turned off (OMC ELOG 209). So it's not totally crazy.)
- [Done] High QE PD preparation / install / QE check
- [Done] High QE PD inventory check
- A1-23 LLO OMC#001
- A1-25 LLO OMC#001
- B1-01 LHO OMC#003
- B1-16 LHO OMC#003
- B1-22 @CIT Cage B1 Cleaned/Installed
- B1-23 @CIT Cage B2 Cleaned/Installed
- C1-03 @CIT Cage B3 Cleaned
- C1-05 Dead / CIT contamination test cav
- C1-07 Dead / CIT contamination test cav
- C1-08 @CIT Cage C2
- C1-09 @CIT Cage C3
- C1-10 @CIT Cage C4
- C1-11 @CIT Cage D1
- C1-12 @CIT Cage D2
- C1-14 @CIT Cage D3
- C1-15 Dead / CIT Cage D4
- C1-17 LHO Spare
- C1-21 LHO Spare
- D1-08 not @CIT, maybe LLO Spare?
- D1-10 not @CIT, maybe LLO Spare?
- [Done] Install & QE check
- [Done] Fiber input beam characterization OMC ELOG 421
== Repair / Preparation ==
- [Done] Obtain: UHV Foil (done) / EP30 (found at the 40m OMC ELOG 434) / EP30-2 bonding setup OMC ELOG 442
- [Done] Crimper tool? LLO bought a new one and modify it.
- [Done] BeCu wire clamps:
- [Done] FirstContact cavity mirror cleaning (see OMC ELOG 414)
- [Done] FirstContact for protection OMC ELOG 430
- Beam dump cleaning -> not necessary
- [Done] Replacing the cable bracket
- [Done] Obtain all parts from Stephen.
- [Done] Class B Torque wrench present in the lab
- [Done] Replacement work
- [Done] Delamination Repair
- [Done] Check all the fasteners / Ready for the shipment OMC ELOG 445
- Check addon mass configuration OMC ELOG 446 OMC ELOG 447
== Shipping ==
- [Done] Tools to ship to LLO OMC ELOG 448
- CLASS B special tool kit (Allens / Pliers / Mighty-Mouse spanner / Spatula / etc)
- FC kit
- Electronic kit (PD connector / trans-impedance amp)
- Spare High QE PDs
- Power meters
- Glass Beamdumps (for optical testing)
- Cable bracket replacement kit (PEEK cable bracket / cable pegs / fastners / spare fasteners / kapton sheet / cable ties)
- Emergency EP30-2 kit (excluding the bond)
- [Done] OMC Pelican Filling (Stephen) / OMC Outerbox/insulation (Stephen/Downs) / OMC Shipment Aug 29, 2022 OMC ELOG 445
|